write the equation of a circle with the center (0,0) and radius 8.
Thank you,
Holly

Answer by John10(297)

write the equation of a circle with the center (0,0) and radius 8.
We know the standard equation for the circle is:
(x - h)^2 + (y - k)^2 = r^2 where (h,k) is the center and r is the radius
Substitute the center (h,k)=(0,0) and r = 8
(x - 0)^2 + (y -0)^2 = 8^2
x^2 + y^2 = 64
Hope it'll help you! Message me if you need any clarification!:)